FOURIER TRANSFORM INFRARED SPECTRA OF Ge_{n} CLUSTERS TRAPPED IN SOLID Ar

\underline{S. DEVDAS},C.M.L. RITTBY, AND W.R.M. GRAHAM, Department of Physics and Astronomy, TCU, Fort Worth, TX 76129.

The structures and vibrational fundamentals of Ge_{n} clusters trapped in solid Ar are under investigation using Fourier transform infrared spectroscopy. Ge_{n}^{−} anions (n=2 - 15) were studied earlier by Neumark et al.^{a}, using anion photoelectron spectroscopy and zero electron kinetic energy spectroscopy. Tentative assignments were made for a fundamental vibration of each of Ge_{3} and Ge_{4}, although vibrational structure was unresolved for larger clusters. We present vibrational spectra obtained for Ge_{n} species produced by laser ablation of pure germanium followed by trapping in an Ar matrix and compare the experimental observations with the predictions of density functional theory calculations. aG. R. Burton, C. Xu, C. Arnold, and D. Neumark, J. Chem. Phys. 104, 2757 (1996).